OpenStudy (anonymous):

solve by elimination method in paired order 6x -9y =7.5 7y -2x = -4.5

OpenStudy (anonymous):

\[6x - {9(2x - 4.5) / 7} = 7.5\] \[6x - (18x + 40.5)/7 = 7.5\] \[42x - 18x + 40.5 = 52.5\] \[24x = 12\] \[x = .5\]

OpenStudy (anonymous):

it is suppose to have a paired order for answer it shows the answer is wrong can u still help me iwth it i put down (12,.5) as answer?

OpenStudy (anonymous):

value of y is -.5. So the paired answer will be (.5, -.5)

OpenStudy (anonymous):

is it ok now?

OpenStudy (anonymous):

6x -9x =19.5 7y -3x =-11

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

is your question right?

OpenStudy (anonymous):

x=5/8 y=-5/12